OpenStudy (anonymous):

thank you what about f+4 and under that g and that =6 for f

OpenStudy (lgbasallote):

who are you thanking?

OpenStudy (lgbasallote):

anyway.. \[\frac{f+4}g = 6\] is that your question?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

start by multiplying g to both sides..what do you get?

OpenStudy (anonymous):

f+4=6f

OpenStudy (anonymous):

hello